OpenStudy (anonymous):

root -48 simplified?

OpenStudy (anonymous):

\[4\sqrt{-3}\]

OpenStudy (austinl):

\[\sqrt{-48}=\sqrt{4}\times\sqrt{-3}\]

OpenStudy (austinl):

\[\sqrt{16}\]correction

OpenStudy (anonymous):

the negative makes it an imaginary number

OpenStudy (austinl):

So you then have \[4i\sqrt{3}\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so do you just simplify it and put the i outside to represent an imaginary number since it's negative?

OpenStudy (austinl):

\[\sqrt{-48}=\sqrt{16}\times\sqrt{-3}\] \[=4\times\sqrt{-3}\] \[=4i\sqrt{3}\]

OpenStudy (austinl):

Yes, essentially!

OpenStudy (anonymous):

okay, thank you!

OpenStudy (austinl):

No problem!